Use the following information to answer questions 6 through 10: Suppose the equation of the line changes to P 5*70. Compute the quantity demanded at each indicated price. 6. Price: $50, Quantity: 7. Price: S40, Quantity: 8. Price: $30, Quantity: 9. Price: $20, Quantity 10. Price: S10 Quantity:
Answer #1

The above question is a simple demand and price equation which can be calculated as follows:-

Question 6

Price=50$

Quantity= Q i.e. to calculate

Equation for this is P= (-5*Q) +70

Post Substitution of values, the equation is as follows:-

50$= (-5*Q) +70

Or

5Q=70-50 i.e. 5Q=20

Therefore Answer to Question 6 is Quantity=20/5=4

Question 7

Price=40$

Quantity= Q i.e. to calculate

Equation for this is P= (-5*Q) +70

Post Substitution of values, the equation is as follows:-

40$= (-5*Q) +70

Or

5Q=70-40 i.e. 5Q=30

Therefore Answer to Question 7 is Quantity=30/5=6

Question 8)

Price=30$

Quantity= Q i.e. to calculate

Equation for this is P= (-5*Q) +70

Post Substitution of values, the equation is as follows:-

30$= (-5*Q) +70

Or

5Q=70-30 i.e. 5Q=40

Therefore Answer to Question 8 is Quantity=40/5=8

Question 9)

Price=20$

Quantity= Q i.e. to calculate

Equation for this is P= (-5*Q) +70

Post Substitution of values, the equation is as follows:-

20$= (-5*Q) +70

Or

5Q=70-20 i.e. 5Q=50

Therefore Answer to Question 9 is Quantity=50/5=10

Question 10)

Price=10$

Quantity= Q i.e. to calculate

Equation for this is P= (-5*Q) +70

Post Substitution of values, the equation is as follows:-

20$= (-5*Q) +70

Or

5Q=70-10 i.e. 5Q=60

Therefore Answer to Question 10 is Quantity=60/5=12

Conclusion:-

This equation is representative of a typical demand schedule, in which as the price for the product goes on declining the quantity demanded for the same increases respectively.

Please feel free to ask your doubts in the comments section